Determine whether the two functions are inverses. f(x) = , and g(x) = 4* © Yes No
Answer #1

Given fw 2 = y(n-3)= 2 lyx - 34-2 2+34 2+37 .: 4-2y) = 2+34 put simplace of y = y + (x) = 2+ 3+ which is equal to g(n) = 2+3x